Skip to content

GE11-P-24L2: Low Optical Power and Odd Current

After modification of the LV boards, the current 1.97A (previously 1.68A) and the optical power decreased.

>>>>>>> RX POWER MEASUREMENT FOR OH #11 <<<<<<<

GBT0 RX power: 89uW
GBT1 RX power: 88uW
GBT2 RX power: 150uW
TRIG0 RX power: 124uW
TRIG1 RX power: 83uW

Previously (after LV filter installation) these were the optical power values:

>>>>>>> RX POWER MEASUREMENT FOR OH #11 <<<<<<<

GBT0 RX power: 363uW
GBT1 RX power: 320uW
GBT2 RX power: 467uW
TRIG0 RX power: 333uW
TRIG1 RX power: 263uW
Edited by Brendan Regnery
To upload designs, you'll need to enable LFS and have an admin enable hashed storage. More information